Fermi National Laboratory


Accelerator Update for October 5, 2005 - October 7, 2005

Wednesday October 5
The day shift began with Operations monitoring store 4431 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, and with MiniBooNE, and the Meson MCenter experiments taking beam. NuMI was not taking beam due to horn problems.

Tevatron (TeV) LCW return pressures came into alarm, experts were called to investigate.

CDF called to notify operations that LOSTP was spiking. After investigation, Tevatron experts were contacted to assist.

At 1041, shot setup to the Recycler for stashing pbars was underway.

The D0 luminosity showed a sudden drop, this was tracked down to a bad card.

Problems with BPM readbacks hinder tuneup, this was resolved.

The beam trajectory out of the Main Injector had to be adjusted and the transfers were about 89% efficient. A -12 volt power supply had to be replaced for the AP3 line BPM's.

Operations notified CDF and D0 that end of store studies would begin at 1500.

Around 1530, the M02 and M03 enclosure interlocks dropped. Experts investigated.

At 1625, the AP3 line BPMs were rebooted.

Pbar studiers turned off the lithium lens for a few minutes while stacking.

At 1813, Store 4431 was officially terminated.

By 1910 four Recycler shots were complete and the Recycler had 282.2E10 pbars spinning.

UPS and QPM work at C4 started at the end of the store and was completed by 1925.

Operators resecured M05, but SY120 beam was disabled for shot setup at 2100.

By 2305, Store 4433 was colliding with an initial luminosity of 140.66e30. The luminosity was a little lower because Recycler beam cooling was more difficult with the larger stash.

Beam returned to Meson MIPP at about 2320.

Thursday October 6
The midnight shift began with Operations monitoring store 4433, with Pbar resuming stacking, and with MiniBooNE, and MCenter taking beam. NuMI is still down for horn troubleshooting and repair.

Tevatron Pbar tune is drifting due to quad roll problems, Operators keeping busy adjusting.

At 0120, MS2LCW system tripped off. Investigation found no power to half the building. FESS called in appropriate people and power was back at 0300.

As fate would have it, Recycler shots began at 0315, so MIPP was not able to start taking beam once things were restored.

By 0415, we were stacking and beam was back to Meson.

At 0805, beam was disabled from Meson for Pbar transfers to the Recycler.

Pbar transfers were done by 0900, beam to MIPP and stacking resumed.

At 0931, problems with a power supply in the Pelletron made an access necessary.

At 0952, BRF7 & 8 tripped on a high vacuum pressure indication.

At 1100, Pbar was allowed to begin reverse proton studies.

Beam headed for switchyard was falling out of the Main Injector, this was tracked down to a spacing between events used by MECAR.

Operators secure MT6 sections 1 & 2 at 1200.

Beam was established to MTest at around 1500.

Meson Center and Test tuning was done around 1740, and both experiments were happy.

MIRF09 tripped several times, but reset after each trip.

Recycler shot preparation began, so beam was turned off to SY120.

Recycler shots began around 1945 and were done by 2030.

At 2235, CHL reports that the A3 house is not looking good. They will try to prevent a quench, but things don't look good.

The store was manually terminated at 2315, due to an unstable fridge.

The A0 Kicker Room air conditioning unit tripped, but FESS restored power.

At 2355, the Tevatron was racked out for an access to hang vacuum pumps on the A37 area.

Friday October 7
The midnight shift began with no store and an access into the Tevatron to add vaccum pumps, with Pbar stacking, monitoring the Recycler stash, and with MiniBooNE taking beam. NuMI is still down with horn work.

At 0030, D0 began their access.

Beam to Meson was off due to the Transfer Hall access.

MIRF9 tripped off a couple of times, but reset each time.

The Plans for Friday and the Weekend
The plans for today are to try to understand the extent of the vacuum problem in the Tevatron. Pbar will stack, Recycler will stash. The NuMI ground fault cleared, but more investigation is required. MiniBooNe, MIPP, and MTest will run over the weekend.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/7/2005   email Fermilab